### v2

# Probing loop quantum effects through solar system experiments: observational signatures and parameter constraints

Abstract
This study investigates quantum gravity effects within the framework of an effective loop quantum gravity black hole (LQG-BH) model parameterized by $\zeta$, utilizing precision measurements from solar system experiments and astrophysical observations. We analyze three classical tests of general relativity (GR): (1) Light deflection constrained by very long baseline interferometry (VLBI) observations of quasar radio signals, (2) Shapiro time delay measurements from the Cassini mission, and (3) Mercury’s perihelion precession determined by MESSENGER mission data. Additionally, we extend our analysis to Earth-orbiting LAGEOS satellites and the relativistic trajectory of the S2 star orbiting the Galactic Center supermassive BH Sagittarius $\rm{A}^{*}$ (Sgr $\rm{A}^{*}$). Our multi-probe approach reveals that the tightest constraint on the LQG parameter comes from Mercury’s perihelion precession, yielding an upper bound $\zeta\lesssim 10^{-2}$. These results establish new observational benchmarks for probing quantum gravity effects.
I Introduction
Over the past century, Einstein’s general relativity (GR) has not only revolutionized our understanding of spacetime and gravity but has also triumphantly survived the most rigorous and precise observational tests across both weak-field and strong-field regimes. In weak-field gravity, GR’s predictions have been validated through precise measurements of astrophysical phenomena such as the perihelion advance of planetary orbits Will:2014kxa ; Park:2017zgd , the deflection of light Fomalont:2009zg , and the Shapiro time delay Bertotti:2003rm . In the strong-field regime, GR has been tested against extreme astrophysical systems, including binary pulsar dynamics Stairs:2003eg ; EventHorizonTelescope:2019dse , black hole (BH) shadow imaging EventHorizonTelescope:2020qrl , and gravitational wave (GW) detections from merging compact objects LIGOScientific:2016aoc . Remarkably, GR’s predictions remain consistent with observations at the current sensitivity levels — a testament to the theory’s enduring robustness and its foundational role in modern physics.
Despite its remarkable theoretical and empirical robustness, GR faces unresolved issues that demand beyond-standard frameworks. These challenges include the theoretical limitations and observational anomalies. Theoretically, GR predicts spacetime singularities at cosmological origins Borde:1993xh and BH centers Hawking:1973uf , where curvature divergences terminate predictability. In addition, no known formalism consistently unifies GR with quantum mechanics Adler:2010wf ; Ng:2003jk , leaving quantum gravity as an open frontier. Observationally, dark matter halos and dark energy, empirically required by $\Lambda$CDM cosmology, lack fundamental justification within GR. Potential tensions in extreme environments, e.g., BH mergers, early-universe physics, may hint at beyond-GR effects.
One of the most effective ways to address these anomalies is to develop a consistent quantum theory of gravity. Among quantum gravity candidates, loop quantum gravity (LQG) provides a non-perturbative, background-independent framework Thiemann:2007pyv ; Smolin:2006pa ; Han:2005km . The cosmological implementation of LQG, known as loop quantum cosmology (LQC), demonstrates singularity resolution by incorporating two key quantum corrections: the inverse volume correction and the holonomy correction Bojowald:2001xe ; Ashtekar:2006rx ; Ashtekar:2006uz ; Ashtekar:2006wn ; Ashtekar:2003hd ; Bojowald:2005epg ; Ashtekar:2011ni . This framework replaces the Big Bang singularity with a nonsingular quantum bounce Ashtekar:2011ni , which then evolves into the current state of the universe Ashtekar:2011ni ; Ashtekar:2015dja . The LQC paradigm naturally extends to spherically symmetric BHs, yielding LQG-BHs. For technical details on LQG-BH construction, see Chiou:2008nm ; Chiou:2008eg ; Boehmer:2007ket ; comprehensive reviews in Perez:2017cmj ; Modesto:2005zm ; Modesto:2008im . In LQG-BHs, the singularity is resolved, and a quantum transition surface typically bridges the trapped and anti-trapped regions Ashtekar:2005qt ; Zhang:2023yps ; Zhang:2023okw ; Zhang:2024ney .
In this paper, we investigate a LQG-inspired BH spacetime featuring double horizons where quantum gravitational effects parameterized by the dimensionless deformation parameter $\zeta$ Zhang:2024khj . While prior studies have extensively characterized this spacetime’s strong-field phenomenology through quasi-normal mode (QNM) spectrum Konoplya:2024lch ; Skvortsova:2024msa ; Zhang:2024svj , photon rings, shadow morphology Liu:2024soc ; Liu:2024wal ; Xamidov:2025oqx , spinning particle dynamics Du:2024ujg , and accretion disk structures Shu:2024tut ; Chen:2025ifv , etc. The BH solution has also been extended to rotating counterparts Ban:2024qsa and scenarios incorporating vacuum charges and cosmological constants Yang:2025ufs . Although prior studies have constrained this model through strong-field observations Konoplya:2024lch ; Heidari:2024bkm ; Wang:2024iwt ; Shu:2024tut ; Chen:2025ifv ; Xamidov:2025oqx , systematic parameter constraints from weak-field regimes, particularly solar system experiments, remain conspicuously absent. We therefore leverage high-precision solar system tests of GR to establish the observational bounds on $\zeta$ [^1] During the final stages of this work, we became aware that constraints on the parameter $\zeta$ have been independently investigated in Xamidov:2025oqx using the MESSENGER mission data from Mercury’s relativistic perihelion precession and the pericenter advance of the S2 star orbiting Sgr $\rm{A}^{*}$..
This paper is organized as follows. Section II provides a thorough geometric characterization of the effective LQG-BH spacetime, with particular emphasis on the geodesic motion of particles in its exterior spacetime. In Section III, the classical GR experiments are employed to probe the effects of the quantum gravity effect. Section IV summarizes the key findings and outlines prospective directions for future investigations. Throughout this work, we adopt geometrized Planck units unless otherwise specified. When using experimental data, the International System of Units (SI) is restored for calculations.
II Test Particle Dynamics in an Effective LQG-BH Spacetime
In this section, we begin by providing a concise overview of the effective LQG-BH model, parameterized by a quantum parameter $\zeta$. We then derive the equations of motion (EOMs) for a test particle orbiting the BH.
II.1 An effective LQG-BH spacetime
The LQG-BH geometry, originally proposed in Zhang:2024khj , is given by:
$\mathrm{d}s^{2}=-{f\left(r\right)}\mathrm{d}t^{2}+\frac{1}{f\left(r\right)}\mathrm{d}r^{2}+r^{2}\left(\mathrm{d}\theta^{2}+\sin^{2}\theta\mathrm{d}\phi^{2}\right)\,,$ (1)
where the metric function $f(r)$ takes the explicit form
$f\left(r\right)=1-\frac{2M}{r}+\frac{M^{2}\zeta^{2}}{r^{2}}\left(1-\frac{2M}{r}\right)^{2}\,.$ (2)
Here, the LQG-corrected parameter $\zeta$ is defined as
$\zeta=\frac{\sqrt{4\sqrt{3}\pi\gamma^{3}\ell_{\mathrm{P}}^{2}}}{M}\,,$ (3)
where $M$ denotes the BH mass, $\gamma$ is the Barbero-Immirzi (BI) parameter of LQG, and $\ell_{\mathrm{P}}$ represents the Planck length. Crucially, since the BI parameter lacks a first-principles determination in current LQG frameworks and is conventionally treated as a free parameter Meissner:2004ju ; Domagala:2004jt , we consequently adopt $\zeta$ as an effective dimensionless free parameter throughout our analysis. This approach allows systematic investigation of LQG-induced modifications to BH.
Figure 1: Left: The metric function $f(r)$ for varying values of the LQG-corrected parameter $\zeta$. The blue curve corresponds to the Schwarzschild case ($\zeta=0$). Right: The inner horizon radius $r_{\mathrm{I}}$ as a function of $\zeta$. The orange line indicates the event horizon $r_{\mathrm{H}}$ for reference.
To analyze the LQG-corrected BH properties, we first examine the metric function $f(r)$ under varying quantum parameter $\zeta$, as illustrated in the left panel of Fig. 1. The equation $g^{rr}=0$, which is equivalent to $f(r)=0$, admits two real roots: one corresponding to the event horizon at $r_{\mathrm{H}}=2M$, and the other representing the inner horizon, given by
$r_{\mathrm{I}}=\frac{M\zeta^{4/3}}{3^{1/3}\left(-9+\sqrt{81+3\thinspace\zeta^{2}}\right)^{1/3}}-\frac{M\zeta^{2/3}\left(-9+\sqrt{81+3\thinspace\zeta^{2}}\right)^{1/3}}{3^{2/3}}\,.$ (4)
The right panel of Fig. 1 illustrates the functional dependence of $r_{\mathrm{I}}$ on $\zeta$. It is evident that in the classical limit of $\zeta$, the spacetime reduces to the Schwarzschild scenario, with the inner horizon $r_{\mathrm{I}}$ collapsing to the singularity at $r=0$. In the finite regime of $\zeta$, i.e., $0<\zeta<\infty$, the inner horizon radius $r_{\mathrm{I}}$ monotonically increases with $\zeta$, maintaining a hierarchy $r_{\mathrm{I}}<r_{\mathrm{H}}$ throughout (see the right panel of Fig. 1). Expanding Eq. (4) in the limit of $\zeta\to\infty$ yields
$r_{\mathrm{I}}=2M-\frac{8M}{\zeta^{2}}+\mathcal{O}\left(\zeta^{-3}\right)\,,$ (5)
demonstrating that $r_{\mathrm{I}}$ asymptotically approaches $r_{\rm{H}}$ while preserving for any finite $\zeta$. This behavior is further corroborated by the left panel of Fig. 1. The introduction of quantum gravity effects resolves the classical Schwarzschild singularity in this spacetime, replacing it with a transition region that connects a BH to a white hole Zhang:2024khj . This region features a bounce surface located within the range $0<r_{\mathrm{B}}<r_{\mathrm{I}}$, where $r_{\mathrm{B}}$ denotes the bounce radius. Such a causal structure aligns closely with those in other LQG-BH models Munch:2021oqn ; Lewandowski:2022zce .
II.2 Test particle dynamics
The Lagrangian governing test particle motion can be expressed as:
$\mathcal{L}\left(x^{\mu},\dot{x}^{\mu}\right)=\frac{1}{2}g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}\,,$ (6)
where the overdot denotes differentiation with respect to the affine parameter $\lambda$ along geodesics. Substituting the spherically symmetric metric (Eq. (1)) into Eq. (6), we obtain the explicit form:
$\mathcal{L}\left(x^{\mu},\dot{x}^{\mu}\right)=\frac{1}{2}\left(-f(r)\dot{t}^{2}+f(r)^{-1}\dot{r}^{2}+r^{2}\dot{\theta}^{2}+r^{2}\sin^{2}\theta\dot{\phi}^{2}\right)\,,$ (7)
The geodesic equations are derived via the Euler-Lagrange formalism:
$\frac{\mathrm{d}}{\mathrm{d}\lambda}\frac{\partial\mathcal{L}}{\partial\dot{x^{\mu}}}-\frac{\partial\mathcal{L}}{\partial x^{\mu}}=0\,.$ (8)
The Euler-Lagrange equation explicitly reveals two conserved quantities associated with spacetime symmetries, namely the energy $E$ and the angular momentum $J$:
$\frac{\partial\mathcal{L}}{\partial\dot{t}}$ $=$ $-f(r)\dot{t}\equiv-E\,,$ (9)
$\frac{\partial\mathcal{L}}{\partial\dot{\phi}}$ $=$ $r^{2}\sin^{2}\theta\dot{\phi}\equiv J\,.$ (10)
These conservation laws originate from the spacetime’s stationarity (time-translation invariance) and axisymmetry (rotational invariance about the polar axis), respectively.
For timelike ($\eta=1$) or null ($\eta=0$) geodesics, the four-velocity satisfies:
$g_{\mu\nu}\frac{dx^{\mu}}{d\lambda}\frac{dx^{\nu}}{d\lambda}=-\eta\,,$ (11)
where $\eta=1$ corresponds to massive particles and $\eta=0$ to massless particles. Constraining the motion to the equatorial plane ($\theta=\frac{\pi}{2}$, $\dot{\theta}=0$) and combining Eq. (11) with the conserved quantities from Eqs. (9) and (10), we derive the radial equation of motion:
$\left(\frac{dr}{d\lambda}\right)^{2}=E^{2}-f\left(r\right)\left(\eta+\frac{J^{2}}{r^{2}}\right)\,,$ (12)
accompanied by the temporal and azimuthal evolution equations:
$\frac{\mathrm{d}t}{\mathrm{d}\lambda}$ $=$ $\frac{{E}}{f\left(r\right)}\,,$ (13)
$\frac{\mathrm{d}\phi}{\mathrm{d}\lambda}$ $=$ $\frac{{J}}{r^{2}}\,.$ (14)
These equations provides the foundation for calculating key observational effects in this LQG-corrected geometry including the light deflection angle, Shapiro time delay, and periastron precession.
III Constraints on quantum parameter
This section derives quantitative constraints on LQG-corrected parameter $\zeta$ through solar system tests. Using the relativistic framework developed in previous sections, we calculate three classical gravitational effects, including the light deflection angle, Shapiro time delay, and periastron precession within the effective LQG-corrected BH spacetime.
III.1 Deflection of light
Consider a light ray propagating along a null geodesic ($\eta=0$) in the solar gravitational field, originating from spatial infinity, reaching a closest approach at radial coordinate $r_{0}$, and escaping back to infinity. The angular deflection per unit radial displacement, derived from the geodesic Eqs. (12) and (14), is governed by
$\frac{\mathrm{d}\phi}{\mathrm{d}r}=\pm\left(\frac{r^{4}}{b^{2}}-f\left(r\right)r^{2}\right)^{-\frac{1}{2}}\,,$ (15)
[…]
At the closest approach $r_{0}$, the radial turning point condition
$\frac{\mathrm{d}r}{\mathrm{d}\phi}\mid_{r=r_{0}}=0\,,$ (16)
directly follows from (15). This imposes a geometric constraint linking $b$ to the spacetime curvature through the metric function $f(r)$:
$b=\sqrt{\frac{r_{0}^{2}}{f\left(r_{0}\right)}}\,.$ (17)
To quantify the cumulative deflection, we compare the total angular change in curved spacetime to the flat-space baseline $\phi=\pi$. The deflection angle $\bigtriangleup\phi$ is thus expressed as
$\bigtriangleup\phi=2\int_{r_{0}}^{\infty}\left(\frac{r^{4}}{b^{2}}-f\left(r\right)r^{2}\right)^{-\frac{1}{2}}\mathrm{d}r-\pi\,,$ (18)
where the factor of $2$ accounts for symmetric deflection during approach and recession. To evaluate the integral in the above equation, we implement the dimensionless substitution $u=\frac{r_{0}}{r}$. In the weak-field regime characterized by $\epsilon\equiv\frac{M}{r_{0}}\ll 1$, we perform a perturbative expansion of $\bigtriangleup\phi$ to second order in $\epsilon$. This yields the asymptotic expression:
$\bigtriangleup\phi$ $=$ $2\int_{0}^{1}\left\{\frac{1}{\sqrt{1-u^{2}}}+\frac{\left(1+u+u^{2}\right)\epsilon}{\left(1+u\right)\sqrt{1-u^{2}}}+\frac{\left[3\left(\frac{1+u+u^{2}}{1+u}\right)^{2}-\left(1+u^{2}\right)\zeta^{2}\right]\epsilon^{2}}{2\sqrt{1-u^{2}}}\right\}\mathrm{d}u$ (19)
-- $\pi\ +\mathcal{O}\left(\epsilon^{3}\right)\,.$
Performing the integration in the above equation, we obtain the leading-order quantum-corrected expression for the light deflection angle:
$\Delta\phi\approx\frac{4M}{r_{0}}\left(1+\frac{15M\pi}{16r_{0}}-\frac{3\pi M\zeta^{2}}{16r_{0}}-\frac{M}{r_{0}}\right)=\Delta\phi_{\rm{GR}}\left[1+\frac{M\left(15\pi-3\pi\zeta^{2}-16M\right)}{16r_{0}}\right]\,,$ (20)
where $\Delta\phi_{\mathrm{GR}}$ represents the standard deflection in GR, with a value of approximately $1.75$ arcsec. We note that the parameter $\epsilon$ has been restored to $\frac{M}{r_{0}}$ in the above expression.
To investigate the detectability of quantum effects and constrain the quantum parameters, we simplify the scenario by defining the closest approach distance $r_{0}$ as the solar radius (corresponding to light grazing the Sun’s limb for detection purposes), while setting $M$ equal to the solar mass. Within the Parameterized Post-Newtonian (PPN) framework, the relativistic gravitational deflection is characterized by the PPN deflection parameter $\gamma$, as illustrated by the expression below:
$\Delta{\phi}_{\rm{PPN}}\approx\Delta{\phi_{\mathrm{GR}}}\left(\frac{1+\gamma}{2}\right)\,.$ (21)
Notice that $\gamma$ strictly equals unity ($\gamma=1$) in GR.
Recent advancements in very long baseline interferometry (VLBI) observations of quasar radio waves deflected by the Sun have yielded unprecedented precision in $\gamma$-determinations Fomalont:2009zg . By integrating upgraded VLBI observational database and advanced analysis frameworks, the deviation of $\left|\gamma-1\right|$ has been improved to the order of $10^{-5}$ lambert2011improved . By incorporating these results and assuming $\zeta>0$ for quantum gravity effect, we compare Eq. (20) with Eq. (21) and directly derive the corresponding bound on $\zeta$ as follows:
$9.12613\displaystyle 0<\zeta<$ 9.12613 . (22)
III.2 Shapiro time delay
The Shapiro time delay, a fundamental gravitational effect predicted by GR, characterizes the increased propagation time of electromagnetic waves as they traverse the curved spacetime in the vicinity of a massive object. This phenomenon has emerged as one of the cornerstone experimental validations of GR. By measuring the PPN parameters, it is possible to investigate the quantum gravity effect and constrain the quantum corrected parameter.
To investigate this, we analyze the superior conjunction in which the satellite and Earth are positioned on opposite sides of the Sun. The radar signals are emitted from Earth, graze the Sun’s gravitational field, and are subsequently reflected by a satellite back to Earth. We begin by deriving the differential equation governing the trajectories of massless particles, expressed in terms of the temporal and radial coordinates $t$ and $r$, through a combination of Eqs. (12) and (13):
$\frac{\mathrm{d}t}{\mathrm{d}r}=\pm\frac{1}{f\left(r\right)\sqrt{1-f\left(r\right)\frac{b^{2}}{r^{2}}}}\,,$ (23)
where the positive and negative signs correspond to the outgoing and incoming trajectories of the radar waves, respectively.
Then, the propagation time of the electromagnetic signal between the closest approach point $r_{0}$ and either the transmitter location $r_{\mathrm{T}}$ on Earth or the satellite receiver location $r_{\mathrm{R}}$ can be formulated as follows:
[…]
where $n=\mathrm{T},\mathrm{R}$. Under weak-field approximation, the propagation time simplifies to:
$\Delta t_{n}$$\approx$ $\sqrt{{r_{n}}^{2}-r_{0}^{2}}+M\left(\sqrt{\frac{r_{n}-r_{0}}{r_{n}+r_{0}}}+2\thinspace\mathrm{arccosh}\left(\frac{r_{n}}{r_{0}}\right)\right)$ (25)
-- $\frac{M^{2}\left[\sqrt{\frac{r_{n}-r_{0}}{r_{n}+r_{0}}}\left(\frac{4r_{n}+5r_{0}}{r_{n}+r_{0}}\right)+6\left(-5+\zeta^{2}\right)\mathrm{arcsin}{\left(\frac{\sqrt{1-\frac{r_{0}}{r_{n}}}}{\sqrt{2}}\right)}\right]}{2r_{0}}\,.$
The leading term $\sqrt{{r_{n}}^{2}-r_{0}^{2}}$ indicates the travel time of radar signals in flat spacetime, whereas the remaining terms encode the additional relativistic time delay corrected by quantum gravity effects. Consequently, the total round-trip time delay for the radar wave propagation can be formally expressed as:
$\Delta t_{\mathrm{SC}}$ $=$ $2\left[\left(\Delta{t}_{\rm{T}}+\Delta{t}_{\rm{R}}\right)-\left(\sqrt{{r_{\rm{T}}}^{2}-r_{0}^{2}}+\sqrt{{r_{\rm{R}}}^{2}-r_{0}^{2}}\right)\right]$ (26)
$\approx$ $4M\left(1+\ln{\left(\frac{4r_{\mathrm{T}}r_{\mathrm{R}}}{r_{0}^{2}}\right)}\right)+\frac{M^{2}\left(15\pi-8-3\pi\zeta^{2}\right)}{r_{0}}\,.$
For comparison, the parametrized PPN gravitational delay per orbit will2018theory ; Weinberg:1972kfs is given by:
$\Delta t_{\rm{PPN}}\approx 4M\left[1+\left(\frac{1+\gamma}{2}\right)\ln\left(\frac{4r_{\mathrm{T}}r_{\mathrm{R}}}{{r_{0}}^{2}}\right)\right].$ (27)
Matching the quantum-gravity-corrected result (Eq. (26)) with the PPN framework (Eq. (27)) yields a direct relation between the PPN parameter $\gamma$ and the quantum gravity corrected parameter $\zeta$:
$\gamma-1=\frac{M\left(15\pi-8-3\pi\zeta^{2}\right)}{2r_{0}\ln\left({\frac{4r_{\mathrm{T}}r_{\mathrm{R}}}{{r_{0}}^{2}}}\right)}.$ (28)
The Cassini solar conjunction experiment, through precision measurements of the Shapiro time delay, currently provides the most stringent observational constraint on the PPN parameter $\gamma$. This yields $\left(\gamma-1\right)=\left(2.1\pm 2.3\right)\times 10^{-5}$ relative to the GR prediction Will:2014kxa ; Bertotti:2003rm . In the actual observations of Cassini’s motion, the heliocentric distances of Earth and the spacecraft were precisely determined as $r_{\mathrm{T}}=1\ \mathrm{AU}$ and $r_{\mathrm{R}}=8.43\ \mathrm{AU}$, respectively, while the radio signal attained its closest solar approach with a radius of $r_{0}=1.6\ R_{\odot}$, where $R_{\odot}$ is the solar radius. Through a systematic comparative analysis, we constrain the quantum-gravity-corrected parameter $\zeta$ to:
$2.60986\displaystyle 0<\zeta<$ 2.60986 . (29)
The Doppler tracking data from the Cassini spacecraft 1992Relativistic ; 1993Doppler provides another approach to constrain the parameter $\zeta$. Unlike direct measurements of the Shapiro time delay, the essence of the Doppler tracking technique lies in measuring the time derivative of Shapiro time delay. Consequently, we obtain the fractional frequency shift of a round-trip radar signal by differentiating Eq. (26) with respect to $t$ Deng:2015sua ; 2000Doppler :
$\delta\nu=\frac{\Delta\nu}{\nu_{0}}=\frac{\mathrm{d}\Delta t_{\mathrm{SC}}}{\mathrm{d}t}\approx\left[-\frac{8M}{r_{0}}-\frac{M^{2}\left(15\pi-8-3\pi\zeta^{2}\right)}{r_{0}^{2}}\right]\frac{\mathrm{d}r_{0}}{\rm{d}t}\approx\delta\nu_{\rm{GR}}+\delta\nu_{\rm{LQG}}\,,$ (30)
where $\Delta\nu\equiv\nu\left(t\right)-\nu_{0}$ quantifies the frequency difference between the Earth-transmitted signal $\nu_{0}$ and the reflected signal received at time $t$. For spacecraft operating at heliocentric distance that significantly exceed Earth’s orbital radius, the time derivative $\mathrm{d}r_{0}/\rm{d}t$ is approximately equivalent to the average orbital velocity of Earth $\nu_{\oplus}$. We then extract the quantum-corrected frequency shift term, which is given as:
$\delta\nu_{\rm{LQG}}\approx\frac{3M^{2}\pi\zeta^{2}}{r_{0}^{2}}\frac{\mathrm{d}r_{0}}{\rm{d}t}=\frac{3M_{\odot}^{2}\pi\zeta^{2}}{R_{\odot}^{2}}\frac{256}{729}\nu_{\oplus},$ (31)
where $M_{\odot}$ denotes the solar mass. Requiring that this quantum correction $\delta\nu_{\rm{LQG}}$ remain below the experimental sensitivity threshold of $10^{-14}$ yields an upper bound on the quantum-gravity-corrected parameter:
$0<\zeta<2.60\,.$ (32)
This constraint exhibits remarkable consistency with bounds derived from Shapiro delay measurements, demonstrating complementary validation through independent relativistic observables.
III.3 Precession of perihelia
In this subsection, we leverage the classical GR prediction of Mercury’s perihelion precession as a precision testbed to quantify quantum gravity effects. Through systematic analysis of orbital dynamics, we establish constraints on the LQG-corrected parameter $\zeta$. To achieve this, we adopt the standard approximation framework in relativistic celestial mechanics by modeling Mercury as a test particle within the Sun’s gravitational field, thereby enabling precise characterization of its geodesic motion.
For timelike geodesics, we have $\eta=1$. It is convenient to adopt the dimensionless inverse radial coordinate $u=\frac{r_{0}}{r}$ as that in Section III.1. Combining Eqs. (12) with (14), the governing differential equation for orbital dynamics takes the following form:
$\left(\frac{\mathrm{d}u}{\mathrm{d}\phi}\right)^{2}=\frac{E^{2}{r_{0}}^{2}}{J^{2}}-f\left(u\right)\left(\frac{{r_{0}}^{2}}{J^{2}}+u^{2}\right)\,.$ (33)
Given the analytical intractability of the exact solution, we employ the perturbative method to solve the above differential equation (33). We begin our perturbative analysis by differentiating Eq. (33) with respect to the azimuthal angle $\phi$. Implementing the weak-field approximation $\epsilon\equiv M/r_{0}\ll 1$, we derive the following expression for relativistic orbital precession:
$\frac{\mathrm{d}^{2}u}{\mathrm{d}\phi^{2}}+u-\frac{M^{2}}{J^{2}{\epsilon}}=-\frac{M^{2}u\zeta^{2}}{J^{2}}+\left(3u^{2}+\frac{6M^{2}u^{2}\zeta^{2}}{J^{2}}\right)\epsilon+\left(-\frac{8M^{2}u^{3}\zeta^{2}}{J^{2}}-2u^{3}\zeta^{2}\right)\epsilon^{2}+\mathcal{O}\left(\epsilon^{3}\right).$ (34)
The LQG imprint emerges crucially through the $\zeta^{2}$-dependent terms.
We implement a recursive perturbative scheme by decomposing the orbital function as $u\left(\phi\right)=u_{0}\left(\phi\right)+u_{1}\left(\phi\right)$ with $u_{0}\left(\phi\right)\ll u_{1}\left(\phi\right)$, where $u_{0}\left(\phi\right)$ represents the dominant Newtonian component and $u_{1}\left(\phi\right)$ encapsulates relativistic-quantum corrections. Truncating at zeroth-order, Eq. (34) admits the unperturbed solution as:
$u_{0}\left(\phi\right)=\frac{M^{2}}{J^{2}\epsilon}\left(1+e\cos\phi\right).$ (35)
The above solution corresponds to the Newtonian Keplerian ellipse parametrized by the classical orbital eccentricity $e$.
To systematically quantify relativistic-quantum corrections, we proceed to determine the first-order perturbation $u_{1}(\phi)$. Implementing the ansatz $u\left(\phi\right)=u_{0}\left(\phi\right)+u_{1}\left(\phi\right)$ with $u_{0}\left(\phi\right)$ given by the Newtonian solution (35), we substitute this decomposition into the precession equation (34). Imposing the boundary conditions $u_{1}\left(0\right)=0$ and ${\mathrm{d}u_{1}\left(0\right)}/{\mathrm{d}\phi}=0$ to ensure continuity with the classical trajectory, the perturbative dynamics are governed by:
$\frac{\mathrm{d}^{2}u_{1}\left(\phi\right)}{\mathrm{d}\phi^{2}}+u_{1}\left(\phi\right)=\sum_{i=0}^{3}\mathcal{P}_{i}\cos^{i}\phi,$ (36)
where the coefficients $\mathcal{P}_{i}$ are given by:
$\mathcal{P}_{0}=\frac{M^{4}\left[3J^{4}-\left(J^{4}-4J^{2}M^{2}+8M^{4}\right)\zeta^{2}\right]}{J^{8}\epsilon},$ (37)
$\mathcal{P}_{1}=\frac{eM^{4}\left[6J^{4}-\left(J^{4}-6J^{2}M^{2}+24M^{4}\right)\zeta^{2}\right]}{J^{8}\epsilon},$ (38)
$\mathcal{P}_{2}=\frac{3M^{4}\thinspace e^{2}\left(J^{4}-8M^{4}\zeta^{2}\right)}{J^{8}\epsilon},$ (39)
$\mathcal{P}_{3}=-\frac{2M^{6}\thinspace e^{3}\zeta^{2}\left(J^{2}+4M^{2}\right)}{J^{8}\epsilon}.$ (40)
Thus, the perturbed part $u_{1}\left(\phi\right)$ is obtained as:
$u_{1}\left(\phi\right)$ $=$ $\mathcal{P}_{0}+\frac{\mathcal{P}_{2}}{2}-\mathcal{P}_{0}\cos\phi-\frac{\mathcal{P}_{2}}{3}\cos\phi+\frac{\mathcal{P}_{3}}{32}\cos\phi-\frac{\mathcal{P}_{2}}{6}\cos\left(2\phi\right)-\frac{\mathcal{P}_{3}}{32}\cos\left(3\phi\right)$ (41)
$+$ $\left(\frac{\mathcal{P}_{1}}{2}+\frac{3\mathcal{P}_{3}}{8}\right)\phi\sin\phi\,.$
Clearly, the relativistic orbital precession behavior of the test particle depends solely on the sine terms. These non-periodic contributions break the azimuthal symmetry—in their absence, the particle would follow closed Keplerian ellipses characteristic of Newtonian gravity. Furthermore, the cumulative effect would render the periapsis deviation observable. The secular accumulation over orbital cycles manifests as observable periastron advance, providing a critical test of spacetime curvature. Retaining only the dominant the sine terms in Eq. (41), the approximate solution to Eq. (34) follows:
$u\left(\phi\right)\approx\frac{Mr_{0}}{J^{2}}\left(1+e\cos\phi\right)+\left(\frac{\chi_{1}}{2}+\frac{3\chi_{3}}{8}\right)\phi\sin\phi\approx\frac{Mr_{0}}{J^{2}}\left\{1+e\cos\left[\left(1-\frac{\delta\phi}{2\pi}\right)\phi\right]\right\}\,,$ (42)
where the cumulative angular precession per orbit
$\delta\phi\approx\frac{6\pi M^{2}}{J^{2}}\left(1-\frac{\zeta^{2}}{6}\right),$ (43)
quantifies the quantum-gravity modified Einstein precession, recovering the classical result when $\zeta\rightarrow 0$. This analytic expression reveals that the LQG correction ($\zeta^{2}$ term) suppress the standard relativistic precession rate.
To establish a direct connection between orbital geometry and relativistic precession with LQG corrections, we leverage the apsidal extremization scheme derived from the radial extrema in Eq. (42). The minimum (pericenter) $r_{-}$ and maximum (apocenter) $r_{+}$ orbital radii occur at occur at angular positions $\left(1-\frac{\delta\phi}{2\pi}\right)\phi=0$ and $\left(1-\frac{\delta\phi}{2\pi}\right)\phi=\pi$ respectively, establishing:
$r_{-}=\frac{J^{2}}{M\left(1+e\right)},$ (44)
$r_{+}=\frac{J^{2}}{M\left(1-e\right)}.$ (45)
From these characteristic radii, we can directly derive the semi-major axis $a$ of any elliptical orbit:
$a=\frac{r_{-}+r_{+}}{2}=\frac{J^{2}}{M(1-e^{2})}.$ (46)
Then, the perihelion advance per orbital revolution is reformulated as:
$\Delta\phi=\frac{6\pi M}{a\left(1-e^{2}\right)}\left(1-\frac{\zeta^{2}}{6}\right)={\Delta\phi_{\mathrm{GR}}}\left(1-\frac{\zeta^{2}}{6}\right),$ (47)
where
${\Delta\phi_{\mathrm{GR}}}=\frac{6\pi M}{a\left(1-e^{2}\right)}.$ (48)
Having established the formalism for relativistic perihelion precession with LQG corrections, we we employ high-precision orbital data from the MESSENGER mission to constrain the LQG-corrected parameter $\zeta$. Our analysis of Mercury’s anomalous perihelion precession yields a precise measurement of $\Delta{\phi}=\left(42.9799\pm 0.0009\right)$ arcsec per century Park:2017zgd , which imposes the following constraint on $\zeta$:
$0.0112089\displaystyle 0<\zeta<$ 0.0112089 . (49)
Furthermore, we extend our analysis to Earth-orbiting LAGEOS satellites and the relativistic trajectory of the S2 star orbiting the Galactic Center supermassive BH Sagittarius $\rm{A}^{*}$ (Sgr $\rm{A}^{*}$). For the LAGEOS satellites, the relativistic perigee precession has been precisely constrained using a 13-year laser-ranging dataset Lucchesi:2010zzb . The observed anomalous precession rate deviates from GR predictions as:
$\Delta\phi=\Delta\phi_{\mathrm{GR}}\left[1+\left(0.28\pm 2.14\right)\times 10^{-3}\right],$ (50)
results in the following bound for $\zeta$:
$0<\zeta<0.105641\,.$ (51)
In the strong-field regime, the first observational test of relativistic periastron advance was performed using the S2 stellar orbit around the galactic center supermassive BH Sgr $\rm{A}^{*}$ Do:2019txf ; GRAVITY:2018ofz ; iorio2024general . By parameterizing deviations through a post-Newtonian inspired parameter $f_{\rm{SP}}$ ($f_{\rm{SP}}=0$ in Newtonian gravity and $f_{\rm{SP}}=1$ in GR), the GRAVITY collaboration analysis GRAVITY:2018ofz establishes a GR-derived periastron shift angle of S2 per orbital period: $\Delta\phi_{\rm{S2}}=\Delta\phi_{\mathrm{GR}}\times f_{\rm{SP}}$, with $\Delta\phi_{\mathrm{GR}}=12.1$ arcmin and $f_{\rm{SP}}=1.1\pm 0.19$. This results in the following constraint:
$0.734847\displaystyle 0<\zeta<$ 0.734847 . (52)
In summary, within the framework of relativistic perihelion precession, the MESSENGER mission currently provides the most stringent constraint on the LQG parameter $\zeta$, surpassing the constraints from Earth-orbiting LAGEOS satellites and the relativistic trajectory of the S2 star orbiting the Galactic Center supermassive BH Sgr $\rm{A}^{*}$. This can be attributed to the higher experimental accuracy of the MESSENGER mission experiment, as noted in GRAVITY:2018ofz ; Chen:2023bao .
IV Conclusion
Table 1: Upper bounds on quantum-gravity-corrected parameter $\zeta$ in the effective LQG black hole spacetime from solar system and galactic center data.
 Experiments/Observations $\zeta$  Datasets
Light deflection 9.126139.126139.12613$\quad$9.12613   VLBI observation of quasars
Shapiro time delay 2.609862.609862.60986$\quad$2.60986   Cassini experiment
2.602.602.60$\quad$2.60  Doppler tracking of Cassini
Perihelion advance 0.01120890.01120890.0112089$\quad$0.0112089  MESSENGER mission
0.1056410.1056410.105641$\quad$0.105641  LAGEOS satellites
0.7348470.7348470.734847$\quad$0.734847   Observation of S2 around Sgr $\rm{A}^{*}$
Classical tests of GR — including light deflection, Shapiro time delay, and perihelion precession — provide fundamental laboratories for probing gravitational theories ranging from GR itself to alternative classical gravities and quantum gravity candidates. In this paper, we investigate quantum gravity effects through an effective LQG-BH model, leveraging precision measurements from these classical GR experiments. Additionally, we also extend our analysis to Earth-orbiting LAGEOS satellites and the relativistic trajectory of the S2 star orbiting the Galactic Center supermassive black hole Sgr $\rm{A}^{*}$. The constraint results are summarized in Table 1.
Theoretical calculations reveal that the inclusion of ${\zeta}^{2}$ terms induces deviations from GR predictions, resulting in a lagged manifestation of relativistic phenomena. As summarized in Table 1, the tightest constraint on $\zeta$ arises from the MESSENGER mission data of the Mercury periasis shift, yielding $0.01120890<\zeta<$ 0.0112089, whereas the second most stringent constraint is produced by the LAGEOS satellites, at the level of $10^{-1}$. To further constrain $\zeta$, we employ the strong gravitational field observations of the S2 star orbit around Sgr $\rm{A}^{*}$, deriving an upper bound $\zeta\lesssim 10^{-1}$, yielding a tighter constraint than those obtained from EHT data of BH shadow radius Konoplya:2024lch ; Heidari:2024bkm ; Wang:2024iwt ; Shu:2024tut ; Chen:2025ifv . Additionally, we estimate the theoretical value of the rescaled parameter $\zeta$ (Eq. (3)) to depend on the solar mass by fixing the BI parameter $\gamma=0.2375$, which gives $\zeta\sim 10^{-39}$. While current solar system experiments lack the resolution to detect LQG signatures, next-generation GW detectors (e.g., LISA LISA:2022yao , Einstein Telescope ET:2019dnz ) may probe such signatures and impose stronger constraints. Furthermore, we anticipate ongoing missions like Gaia Gaia:2016zol and BepiColombo Will:2018mcj , along with proposed space-based projects such as LATOR LATOR:2005hia and BEACON Turyshev:2008rh , will progressively improve the precision of LQG-corrected parameters constraints.
Future investigations could extend this research program along the following promising avenues. First, the methodology developed herein may be extended to analyze quantum corrections in gravitational time delay effects generated by spinning oblate masses as Battista:2017xlm . Second, a particularly relevant application would involve re-examining the Newtonian Lagrangian points within the Earth-Moon system through the lens of quantum gravity phenomenology, potentially revealing observable signatures in celestial mechanics Battista:2014oba ; Battista:2014ija ; Battista:2015zta ; Battista:2015wwa ; Tartaglia:2018bjc ; Battista:2020qqp . In addition, the emerging framework of LQG could be rigorously tested through detailed simulations of extreme mass-ratio inspirals, where recent theoretical advances Fu:2024cfk ; Zi:2024jla suggest new observational windows for probing quantum spacetime structure.
Acknowledgements.
We are especially grateful to Prof. Rui-Hong Yue for helpful discussions and suggestions. This work is supported by the Natural Science Foundation of China under Grants No. 12375055.